OpenStudy (anonymous):

Simplify Completely 8x^3+6x^2-4x over -2x

OpenStudy (anonymous):

\[8x^3+6x^2-4x \over -2x \] Completely Simplify

OpenStudy (diyadiya):

divide each term by -2x

OpenStudy (anonymous):

each term on top? or all of them?

OpenStudy (diyadiya):

\[\frac{8x^3}{-2x} + \frac{6x^2}{-2x}- \frac{4x}{-2x}\]can you do that?

OpenStudy (anonymous):

i just did..now im about to get the answer.

OpenStudy (anonymous):

i think...i hope lol

OpenStudy (diyadiya):

what did you get?

OpenStudy (anonymous):

I got -4x^2-3x-2x^0...?

OpenStudy (diyadiya):

-4x^2-3x-(-2x^0) x^0 =1 so -4x^2-3x + 2

OpenStudy (anonymous):

Yea thats what I got :)

OpenStudy (diyadiya):

Good :)

OpenStudy (anonymous):

thank you so much...your the best!! :)

OpenStudy (diyadiya):

You're most welcome :)